if (1/ab)+(1/bc)+(1/ac)=1 and a+b+c=1 find value of (1/(1+a+ab))+(1+(1+b+bc))+(1+(1+c+ac))

Grade:9

1 Answers

DASARI NAGESWARA RAO
6 years ago
Given that 1/ab+1/bc+1/ca =1 and A+B+c=1.1/ab+1/bc+1/ca=1=>(a+b+c)/abc =1 abc=1=>bc=1/a,c=1/ab,ac=1/b.(1/1+a+ab)+(1/1+b+bc)+(1/1+c+ac)=(1/1+a+ab)+(1/1+b+1/a)+(1/1+1/ab+1/b)=(1/1+a+ab)+(a/1+a+ab)+(ab/1+a+ab)=(1+a+ab)/(1+a+ab)=1